OpenStudy (noahswager):

how do i solve square root of 6x+ 7 square root of 216x- 6 square root of 150x?

OpenStudy (tayfishb):

\[\sqrt{6x+7} \sqrt{216x-6} \sqrt{150x} ?\]

OpenStudy (noahswager):

no the +7 is in front of the square root of 216x and same with the -6

OpenStudy (noahswager):

i may have figured out how to do it though

OpenStudy (tayfishb):

ok well you know the rule thats like \[\sqrt{a} \sqrt{b} \sqrt{c}=\sqrt{abc}\] right?

OpenStudy (noahswager):

yea i got the answer thank you